OpenStudy (anonymous):

translate this word problem as a system of equation using subtraction find to numbers who sum is 58 and who difference is 16

OpenStudy (anonymous):

can any one help pls

OpenStudy (triciaal):

find TWO numbers sum = + difference = - let the numbers be a and b a + b = 58 a - b = 16 to solve add the two eqn 2a = 74 divide by 2 a = 37 now substitute in one of the eqn a + b = 58 37 + b = 58 so b = 58-37 = 21 to check your answer a - b = 16 37 - 21 = 16 as stated